OpenStudy (anonymous):

The equation of Circle P is ( x − 3)2 + (y + 2)2 = 34 and the equation of circle Q is (x − 4)2 + (y + 5)2 = 52. What is the distance between the centers of these circles?

OpenStudy (anonymous):

Please help... I'd prefer if someone would actually help me how to solve it rather then just giving me an answer.

OpenStudy (hoblos):

the center of p is O=(3,-2) the center of Q is O'=(4,-5) d=OO' = sqrt(10)

OpenStudy (anonymous):

So the distance ~3.16 ?